Question 11

1 Point

The Ksp of Ag2CO3 is 8.11 x 10-12. What is the maximum Ag+ ion concentration possible if the carbonate ion concentration is 1.14 x 10-4 M?

Select the correct answer

  1. 1.14 x 10-4 M

  2. 2.67 x 10-4 M

  3. 2.24 x 10-4 M

  4. 7.11 x 10-8 M
Answer #1

the Ag2CO3 will dissociates as

Ag2CO3\rightleftharpoons 2Ag+ + CO32-  

and the Ksp for this reaction is

Ksp = [Ag+]2[CO32-]

Now we have

Ksp = 8.11 \times 10-12

[CO32-] = 1.14 \times 10-4 M

hence by using above Ksp formula we get

Ksp = [Ag+]2[CO32-]

8.11 \times 10-12 = [Ag+]2\times 1.14 \times 10-4

[Ag+]2 = 8.11 \times 10-12 / 1.14 \times 10-4

= 7.11 \times 10-8

by taking square roots on bothe sides we get

[Ag+] = 2.67 \times 10-4 M

so our answer is

[Ag+] = 2.67 \times 10-4 M